# DMSC: Dynamic Multi-Scale Coordination Framework for Time Series Forecasting

### v2

@@ Introduction
To address the limitations of static fusion mechanisms in multiscale forecasting, we propose the Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) block. This dynamic prediction head models temporal dependencies at different horizons via hierarchical experts (global-shared and local-specialized experts) and adaptively weights scale-specific predictions based on temporal patterns.
Specifically, ASR-MoE comprises two expert groups to handle distinct temporal granularities. Global experts $\mathcal{E}^{g}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{m}\}$ capture common long-term dependencies through deep nonlinear transformations, modeling persistent trends and long-periodic patterns that commonly existed in time series. Local experts $\mathcal{E}^{l}=\{\mathcal{L}_{1},\mathcal{L}_{2},...,\mathcal{L}_{n}\}$ specialize in different short-term variations through shallower networks, detecting short-periodic patterns and high-frequency fluctuations that dynamically varied across stages. This explicit decoupling enables specialized handling of complex dynamics while maintaining higher parameter efficiency.
For weights calculation of each scale’s experts, a dynamic routing mechanism is proposed to assign input-dependent weights to experts, balancing global and local dependencies. Given a scale-specific feature $\mathbf{F}_{l}$, the router generates weights as follows:
@@
where the weight matrix $\mathbf{\Omega}$ are divided into global weights $\mathbf{\Omega}^{G}$ and local weights $\mathbf{\Omega}^{L}$. For computational efficiency, we activate only the top-$K$ local experts via sparse routing:
$\mathbf{\hat{\Omega}}^{L}_{l},\mathcal{I}=\mathrm{Top-K}(\mathbf{\Omega}^{L}_{l},K),$ (11)
where $\mathcal{I}$ denotes indices of selected experts.
Finally, a temporal-aware weighting module fuses outputs from all scales by historical scale importance. Scale-specific predictions $\hat{\mathbf{Y}}^{l}$ are fused using time-dependent weights:
$\mathbf{w}=\mathrm{Softmax}(\mathrm{MLP}(\bigoplus^{L}_{l=1}\phi(\mathbf{F}_{l})\cdot\mathbf{w}_{hist})),$ (12)
where $\phi(\cdot)$ denotes temporal descriptors, $\mathbf{w}_{hist}$ denotes historical weighting memory, and $\bigoplus$ means concatenation. The final prediction integrates multi-scale outputs:
@@
Meanwhile, to ensure balanced expert utilization, ASR-MoE incorporates a balance loss:
@@
and the overall loss function of DMSC is as follows:
@@
By combining hierarchical expert specialization, dynamic routing, and temporal-aware weighting, ASR-MoE enables the proposed framework to adaptively prioritize relevant scales and experts, achieving a synergy between long-term trend projection and short-term detail refinement in TSF.
@@
To comprehensively verify the performance and effectiveness of DMSC, we conduct extensive experiments on 13 real-world TSF benchmarks spanning multiple domains and temporal resolutions.
@@
For long-term forecasting, we conduct experiments on nine well-established benchmarks, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, ECL, Traffic, Weather and Solar. For short-term forecasting, we experiment on PEMS(PEMS03, PEMS04, PEMS07, PEMS08).
@@
The main results are listed in Table 1 for long-term forecasting and Table 2 for short-term forecasting, in which the lower MSE and MAE indicates better forecasting performance. Our proposed DMSC framework demonstrates consistent SOTA performance across all 13 benchmarks, achieving the lowest MAE and MSE in most experimental configurations. These results confirm robust effectiveness and generality of DMSC for both long-term and short-term forecasting tasks. Notably, multi-scale architectures like TimesNet and PatchMLP exhibit instability due to their inflexible fusion mechanisms that fails to fully leverage captured multiscale features. And the preferable performance of TimeXer and iTransformer underscores the effect of the patch-wise and variate-wise strategy in information representation, but their fixed decomposition schemes inherently constrain the model’s potential. Compared to iTransformer and Dlinear, DMSC has a considerable improvement in data with more variate(Traffic, PEMS), demonstrating the necessity of jointly dependency modeling. Collectively, the above observations validate DMSC’s efficacy in addressing the core challenges of time series forecasting.
@@
For EMPD, employing predefined static patches or raw data embedding instead of input-adaptive decomposition degraded performance. This demonstrates the necessity of patch-wise decomposition and lightweight neural controllers for granularity adjustment. For TIB, replacing joint modeling with standard convolution layers or only intra-patch features resulted in performance drops, validating the need for comprehensive and joint modeling of triadic dependencies. Meanwhile, ablation of dynamic fusion mechanism in TIB shows its effectiveness to dynamically fuse these heterogeneous dependencies, as static handling of such relationships undermines holistic dependency capture across different scales. For ASR-MoE, substituting our adaptive prediction mechanism with either a single prediction head or summation of multiple linear heads yields substantial performance declines, which demonstrates that single heads inadequately utilize multiscale features, and simple aggregation of multiple heads overlooks the differential importance of scale-specific dependencies. Moreover, removing either global or local experts leads to performance degradation, which demonstrates the necessity of leveraging distinct experts for different temporal patterns. The elaborate design in ASR-MoE for collective utilization of specialized experts can effectively harness diverse temporal dynamics.
Collectively, these ablation studies confirm that dynamic and multiscale modeling is central to DMSC. All of three blocks and the proposed progressive cascade architecture together enable robust multiscale modeling and hierarchical feature learning. The consistent performance drops from ablating any component validate the integrity of DMSC’s architecture.
@@
To verify the specialization of ASR-MoE experts, we conducted activation pattern visualization across different experts. As shown in Fig. 2, distinct specialization profiles emerge: global experts exhibit pronounced sensitivity to persistent trends and long-periodic patterns, while local experts demonstrate acute responsiveness to short-periodic patterns and high-frequency fluctuations. This specialization demonstrates adaptively knowledge learning for different experts.
@@
Theoretically, longer input sequences provide richer historical information, potentially enhancing the model’s ability to capture long-term trends and complex multiscal dependencies. Our DMSC framework is specifically designed to leverage this, with its dynamic multiscale coordination capability inherently enabling adaptive modeling and effective utilization of information across varying look-back lengths. So we evaluate DMSC with look-back length from {48, 96, 192, 336, 720} on Electricity dataset. As shown in Fig. 3, DMSC consistently improves performance with longer historical sequences while outperforms other baselines, demonstrating its capacity to extract richer temporal representations from extended contexts and effectiveness in integrating multiscale temporal patterns.
@@
We rigorously evaluate the memory usage and training time of DMSC against other SOTA baselines on Weather(21 variates) and Traffic datasets(862 variates). As demonstrated in Fig. 4 and Fig. 5, DMSC consistently achieves superior efficiency in both memory usage and training time while maintaining competitive forecasting accuracy. Crucially, DMSC demonstrates near-linear scalability to long sequences—significantly outperforming Transformer-based models that exhibit quadratic complexity and MLP-based approaches with parameter explosion. These efficiency gains are particularly pronounced when handling high-dimensional multivariate data (Traffic), where DMSC’s cross-variable interaction avoids the computational overhead of exhaustive pairwise attention while maintaining modeling capacity. Our framework thus establishes a balance of accuracy, latency, and memory usage for practical deployment scenarios.
@@ Conclusion
This paper introduces the Dynamic Multi-Scale Interaction (DMSC) framework, a novel framework that can advance TSF through dynamic multiscale coordination. DMSC has three key contributions: 1) an Embedded Multi-Scale Patch Decomposition(EMPD) is introduced to dynamically decompose time series into hierarchical patches to eliminate predefined and fixed scale constraints; 2) Triad Interaction Block (TIB) is proposed to jointly model intra-patch, inter-patch, and cross-variable dependencies, thus forming a coarse-to-fine feature pyramid through progressive cascade layers; 3) an Adaptive Scale Routing MoE (ASR-MoE) is proposed to dynamically fuse multiscale predictions via temporal-aware weighting of global and local experts. Extensive experiments on thirteen benchmarks demonstrate that DMSC can achieve SOTA performance and superior efficiency. Future work will extend the framework to multi-tasks learning and explore optimizations on intricate dataset environments, improving its multi-scenario applicability for real world.
@@ Appendix B Model Analysis
The embedding dimension determines the richness of feature representations. We set $d_{\mathrm{model}}\in\{64,128,256,512\}$ to evaluate model’s performance on ECL and Weather dataset. The results are shown in Fig. 6(a) and Fig. 6(b), we choose the optimal dimension at 128 / 256 - smaller embedding impairs multiscale feature separation, while larger configuration yield marginal gains with higher memory cost, indicating dimensionality saturating.
@@
EMPD employs exponentially decaying patch lengths to capture multiscale dependencies. We vary the decay rate (2, 3, 4) and find that a decay factor of 2 yields the best results from Fig. 7(a) on weather dataset. It can be seen that a decay factor of 2 yields the best forecasting performance, which balances the preservation of sufficient local details in fine-grained layers against the extraction of broader trends in coarse-grained layers.
@@
The number of cascade layers ($l\in\{1-5\}$) affects the hierarchical feature extraction. As shown in Fig. 7(b), three layers can achieve peak efficiency - shallower stacks (1 - 2 layers) preserve perfect accuracy via residual propagation but miss fine-grained interactions, while deeper configurations (4 - 5 layers) introduce diminishing returns with much more latency but fewer improvement.

### v4

@@ Introduction
To address the limitations of static fusion mechanisms in multi-scale forecasting, we propose the Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) block. This dynamic prediction head captures temporal dependencies at different horizons through a hierarchy of experts (global-shared and local-specialized experts) and adaptively aggregates scale-specific predictions based on temporal patterns.
Specifically, ASR-MoE incorporates two distinct expert groups to handle different temporal granularities. Global experts $\mathcal{E}^{g}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{m}\}$ capture common long-term dependencies through deeper networks, modeling persistent trends and long-periodic patterns that commonly existed in time series. Local experts $\mathcal{E}^{l}=\{\mathcal{L}_{1},\mathcal{L}_{2},...,\mathcal{L}_{n}\}$ capture diverse short-term variations via shallower networks, detecting short-periodic patterns and high-frequency fluctuations that vary dynamically across stages. This explicit decoupling enables specialized handling of complex temporal dynamics while maintaining high parameter efficiency.
To compute the weighting for experts at each scale, a dynamic routing mechanism assigns input-dependent weights, thereby balancing the contributions of global and local dependencies. Given a scale-specific feature $\mathbf{F}_{l}$, the router generates weights, and only the top-$K$ local experts are activated via sparse routing:
$\mathbf{\hat{\Omega}}^{L}_{l},\mathcal{I}=\begin{cases}\mathbf{\Omega}^{L}_{l},&\mathbf{\Omega}^{L}_{l}\in\mathrm{Top-K}(\mathbf{\Omega}_{l},K)\\ 0,&\text{otherwise}\end{cases}$
$\mathbf{\Omega}_{G}=\mathrm{Sigmoid}(\mathrm{MLP}(\mathbf{F}_{l})),$
@@
where $\mathbf{\Omega}_{g}$ and $\mathbf{\Omega}_{l}$ represent the global and local expert weight matrices, respectively. $\mathcal{I}$ denotes the set of indices for the selected experts. Finally, a temporal-aware weighting module fuses outputs from all scales by historical scale importance. The scale-specific predictions $\hat{\mathbf{Y}}^{l}$ are fused using time-dependent weights:
$\mathbf{w}=\mathrm{Softmax}\left(\mathrm{MLP}\left(\bigoplus^{L}_{l=1}\phi(\mathbf{F}_{l})\cdot\mathbf{w}_{\text{hist}}\right)\right),$ (11)
where $\phi(\cdot)$ denotes temporal descriptors, $\mathbf{w}_{\text{hist}}$ represents the historical weighting memory, and $\bigoplus$ denotes the concatenation operation. The final prediction is obtained by integrating the multi-scale outputs:
@@
To ensure balanced expert utilization, ASR-MoE incorporates an auxiliary balance loss:
@@
The overall loss function for DMSC is defined as follows:
@@
By integrating hierarchical expert specialization, dynamic routing, and temporal-aware weighting, ASR-MoE adaptively prioritizes relevant scales and experts, achieving a synergistic balance between long-term trend capture and short-term detail refinement in time series forecasting.
@@
To comprehensively evaluate the performance and effectiveness of the proposed DMSC framework, we conduct extensive experiments on 13 real-world TSF benchmarks across multiple domains and temporal resolutions.
@@
For long-term forecasting, we conduct experiments on nine well-established benchmarks, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, ECL, Traffic, Weather and Solar. For short-term forecasting, we utilize the PEMS(PEMS03, PEMS04, PEMS07, PEMS08) dataset.
@@
The main results for long-term and short-term forecasting are presented in Table 1 and Table 2, in which the lower MSE and MAE values indicates superior forecasting performance. The proposed DMSC framework demonstrates consistent SOTA performance across all 13 benchmarks, achieving the lowest MAE and MSE in most experimental settings. These results confirm the robust effectiveness and generalizability of DMSC for both long-term and short-term forecasting tasks. Notably, multi-scale architectures like TimesNet and PatchMLP exhibit performance instability, which we attribute to their inflexible fusion mechanisms that fails to fully leverage the captured multi-scale features. The competitive performance of TimeXer and iTransformer underscores the effectiveness of patch-wise and variate-wise strategy for information representation, but their fixed decomposition schemes inherently constrain the model’s potential. Furthermore, compared to iTransformer and Dlinear, DMSC achieves considerable improvement in data with more variate(Traffic, PEMS), highlighting the necessity of jointly dependency modeling. Collectively, the above observations validate the efficacy of DMSC in addressing the core challenges of time series forecasting.
@@
For EMPD, replacing the input-adaptive decomposition with either predefined static patches or raw data embedding leads to performance degradation. This demonstrates the necessity of adaptive patch-wise decomposition and its lightweight network for dynamic granularity adjustment. For TIB, replacing joint modeling with standard convolution layers or only intra-patch features resulted in performance drops, validating the necessity of comprehensive joint modeling of triadic dependencies. Similarly, ablation of dynamic fusion mechanism in TIB confirms its effectiveness in fusing heterogeneous dependencies, as static handling undermines holistic cross-scale dependency capture. For ASR-MoE, replacing the adaptive prediction mechanism with a single prediction head or summed multiple linear projections leads to substantial performance declines. This indicates that single heads inadequately utilize multi-scale features, while simplistic aggregation overlooks scale-specific dependency importance. Moreover, removing either global or local experts degrades performance, demonstrating the necessity of maintaining distinct experts for different temporal patterns. The elaborate design of ASR-MoE effectively harnesses diverse temporal dynamics through collective specialized expert utilization. Collectively, these ablation studies confirm that dynamic and multi-scale modeling is central to DMSC. All of three blocks and the proposed progressive cascade architecture together enable robust multiscale modeling and hierarchical feature learning, with consistent performance drops from ablating any component validate architectural integrity of DMSC.
@@
To verify the specialization of ASR-MoE experts, we visualize their activation patterns across different temporal experts. As shown in Fig. 2, distinct specialization profiles emerge: global experts exhibit heightened sensitivity to persistent trends and long-periodic patterns, whereas local experts demonstrate acute responsiveness to short-periodic patterns and high-frequency fluctuations. This observed specialization demonstrates the adaptive acquisition of specialized knowledge by different experts.
@@
Theoretically, longer input sequences provide richer historical information, which can enhance the model’s ability to capture long-term trends and complex multi-scale dependencies. Our DMSC framework is specifically designed to leverage this characteristic, as its dynamic multi-scale coordination capability inherently facilitates adaptive modeling and effective information utilization across varying look-back lengths. So we evaluate DMSC with look-back lengths selected from {48, 96, 192, 336, 720} on Electricity dataset. As shown in Fig. 3, DMSC consistently achieves improved performance with longer historical sequences while outperforming other baseline methods. This demonstrates its capacity to extract richer temporal representations from extended contexts and effectively integrate multi-scale temporal patterns.
@@
We rigorously evaluate the memory usage and training time of DMSC against other SOTA baselines on Weather(21 variates) and Traffic(862 variates) datasets. As demonstrated in Fig. 4 and Fig. 5, DMSC consistently achieves superior efficiency in both memory usage and training time while maintaining competitive forecasting accuracy. Crucially, DMSC demonstrates near-linear scalability to long sequences, as outperforming Transformer-based models with quadratic complexity and MLP-based approaches susceptible to parameter explosion. These efficiency gains are particularly pronounced when processing high-dimensional multivariate data (Traffic), where the cross-variable interaction in TIB avoids the computational overhead of exhaustive pairwise attention while preserving modeling capacity. DMSC thus establishes a balance of accuracy, latency, and memory usage for practical deployment scenarios.
@@ Conclusion
This paper presents the Dynamic Multi-Scale Coordination (DMSC) framework, a novel approach that advances time series forecasting through comprehensive dynamic multi-scale coordination. DMSC has three key contributions: 1) Embedded Multi-Scale Patch Decomposition(EMPD) dynamically decomposes time series into hierarchical patches, eliminating predefined scale constraints through input-adaptive granularity adjustment; 2) Triad Interaction Block (TIB) jointly models intra-patch, inter-patch, and cross-variable dependencies, forming a coarse-to-fine feature pyramid through progressive cascade layers; 3) Adaptive Scale Routing MoE (ASR-MoE) dynamically fuses multi-scale predictions via temporal-aware weighting of specialized global and local experts. Extensive experiments on thirteen benchmarks demonstrate that DMSC achieves SOTA performance and superior efficiency. Future work will extend the framework to multi-task learning and explore optimizations on complex real-world environments, thereby enhancing its applicability across diverse scenarios.
@@ Appendix B Model Analysis
The embedding dimension determines the richness of feature representations. We evaluate the model’s performance on ECL and Weather datasets with embedding dimensions $d_{\mathrm{model}}\in\{64,128,$ 256,512}256,512\}. The results in Fig. 6(a) and Fig. 6(b) show that the optimal dimension is at 128 / 256, as smaller embedding impairs multi-scale feature separation, while larger configurations yield marginal gains with higher memory cost, indicating dimensionality saturating.
@@
EMPD employs exponentially decaying patch lengths across layers to capture multi-scale dependencies. We evaluate the impact of the EMPD decay rate by testing values of $\tau\in\{2,3,4\}$ on Weather dataset. As shown in Fig. 7(a), a decay factor of $\tau=2$ achieves the best forecasting performance, which balances the preservation of sufficient local details in fine-grained layers with the extraction of broader trends in coarse-grained layers.
@@
The number of progressive cascade layers ($l\in\{1,2,3,4,5\}$) influences hierarchical feature extraction. As shown in Fig. 7(b), three layers can achieve optimal efficiency, shallower stacks (1 - 2 layers) fail to capture fine-grained interactions, while deeper configurations (4 - 5 layers) introduce substantially increased latency with diminishing performance returns.